# Spirals, rings, and vortices shaped by shadows in protoplanetary disks: from radiative hydrodynamical simulations to observable signatures

### v1

@@ Appendix C Azimuthal $T$ variations in RADMC-3D
As mentioned in Sect. 5.1, RADMC-3D is a hydrostatic code and therefore does not capture the azimuthal advection of thermal energy. This results in artificially narrower temperature fluctuations compared to our hydrodynamical simulations. To highlight this effect we show the azimuthal deviations of $T_{\text{small}}\approx T_{\text{gas}}$ from the RADMC-3D output of our fiducial model at $t=1000\,P_{0}$ and $R=30$ au in Fig. 20 and compare it to the PLUTO output at the same timestamp and radius.
Figure 20: Azimuthal slice at $R=30$ au of the gas temperature deviation, normalized to an azimuthal average for RADMC-3D (blue, orange) and to the initial temperature at that radius for PLUTO (black). The latter shows a wider spread of temperatures due to the azimuthal advection of thermal energy.

### v2

@@ Appendix C Azimuthal $T$ variations in RADMC-3D
As mentioned in Sect. 5.1, RADMC-3D is a hydrostatic code and therefore does not capture the azimuthal advection of thermal energy. This results in artificially narrower temperature fluctuations compared to our hydrodynamical simulations. To highlight this effect we show the azimuthal profiles of $T_{\text{small}}\approx T_{\text{gas}}$ from the RADMC-3D output of our fiducial model at $t=1000\,P_{0}$ and $R=40$ au in Fig. 20 and compare it to the PLUTO output at the same timestamp and radius.
As seen on that figure, deviations from said background are found only around the shadowed region in RADMC-3D whereas they span the full azimuthal range in PLUTO. The RADMC-3D model also shows weaker deviations compared to the background near the midplane compared to PLUTO. Nevertheless, the two codes agree very well at the midplane in terms of the background temperature.
Figure 20: Azimuthal slice at $R=40$ au of the gas and small dust temperatures, normalized to the azimuthal median for RADMC-3D (blue, orange) and to the initial temperature at that radius for PLUTO (black). The latter shows a wider spread of temperatures due to the azimuthal advection of thermal energy. RADMC-3D and PLUTO show excellent agreement in background temperature at the midplane.